Honda:

Can someone explain what density is and how to find it? I'm using this for geometry and have no clue how to do it.

Honda:

here's an example of what i'm talking about if that helps.

KiWiPoP1:

Density determines how compact a substance is, for this you need to use an equation which is p=m/v

KiWiPoP1:

so you need to find your volume and your mass

KiWiPoP1:

52 would be your volume, and 19 would be your mass, I'm not too sure, maybe someone can back me up on this

KiWiPoP1:

p= m/v which in this case is also p=19/54

KiWiPoP1:

p=density

Honda:

okay, thank you!

KiWiPoP1:

NP~